ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 233, 2019

Belov, A. and A. Erschler No iterated identities satisfied by all finite groups

167

Bernazzani, D. Centralizers in the group of interval exchange transformations

29

Betta, M. F., F. Feo and M. R. Posteraro Uniqueness results for strongly monotone operators related to Gauss measure

297

Billig, Y., V. Futorny and J. Nilsson Representations of Lie algebras of vector fields on affine varieties

379

Bownik, M. and I. Londner On syndetic Riesz sequences  113
Brian, W. R. Universal flows and automorphisms of P(ω)/fin  453

Castillo-Ramirez, A. and J. McInroy Code algebras which are axial algebras and their ℤ_{2}-gradings

401

Conlon, D. Hypergraph expanders from Cayley graphs  49

Conlon, D., J. Fox, M. Kwan and B. Sudakov Hypergraph cuts above the average

67

Debruyne, G. and D. Seifert An abstract approach to optimal decay of functions and operator semigroups

439

Deng, H., H. Liu and L. Tian Critical points of solutions for the mean curvature equation in strictly convex and nonconvex domains

311

Eleftheriou, P. E. Small sets in dense pairs  1

Erschler, A. See A. Belov

Eskew, M. More rigid ideals  225

Feo, F. See M. F. Betta Fox, J. See D. Conlon Futorny, V. See Y. Billig Herzog, J., T. Hibi and D. I. Stamate The trace of the canonical

module

133

Hibi, T. See J. Herzog Kwan, M. See D. Conlon Lankeit, J. and M. Winkler Counterintuitive dependence of tem-

poral asymptotics on initial decay in a nonlocal degenerate parabolic equation arising in game theory

249

Liu, H. See H. Deng Londner, I. See M. Bownik McInroy, J. See A. Castillo-Ramirez Nilsson, J. See Y. Billig Posteraro, M. R. See M. F. Betta Seifert, D. See G. Debruyne Soukup, D. T. A model with Suslin trees but no minimal uncountable

linear orders other than ω_{1} and −ω_{1}

199

Stamate, D. I. See J. Herzog Sudakov, B. See D. Conlon Tian, L. See H. Deng

Varjú, P. P. Expansion of coset graphs of PSL_{2}(F_{p})  335

Wilkes, G. Profinite rigidity of graph manifolds, II: knots and mapping classes

351

Winkler, M. See J. Lankeit